PLEASE HELP SOLVE
FOR EACH PAIR OF FUNCTIONS, FIND (A)(F+G)(X) AND (B) (F-G)(X)
SOLVE: F(X)=5X-10, G(X)=3X+7

Found 2 solutions by jai_kos, praseena:
Answer by jai_kos(139)

a) (F + G) (x) = F(x) + G(x) = 5x – 10 + 3x + 7
= 5x + 3x -10 + 7
= 8x - 3
b) (F- G) (x) = F(x) – G(x) = 5x -10 – (3x + 7)
= 5x -10 – 3x – 7
= 5x -3x -10 -7
= 2x – 17

Answer by praseena(37)

FOR EACH PAIR OF FUNCTIONS, FIND (A)(F+G)(X) AND (B) (F-G)(X)
SOLVE: F(X)=5X-10, G(X)=3X+7
---------------------------------------------------------------------------
SOLUTION:

Let f and g be two functions. The sum, f + g, the difference, f – g, the product, fg, and the quotient, are functions defined as follows:
1. Sum: (f + g)(x) = f(x) + g(x)
2. Difference: (f – g)(x) = f(x) – g(x)
-----------------------------------------------------------------
Let's come back to your solutions:
FIND (A)(F+G)(X) AND (B) (F-G)(X)

F(X)=5X-10,

G(X)=3X+7


(F+G) x= F(x) + G(x)
= 5x -10 + 3x +7
= 5x + 3x -10 +7
= 8x - 3
(F-G)x = F(x) - G(x)
= 5x -10 - (3x+7)
= 5x -10 -3x - 7
= 5x - 3x -10 -7
= 2x -17